### v1

# AlignAR: Generative Sentence Alignment for Arabic–English Parallel Corpora of Legal and Literary Texts

Abstract
High-quality parallel corpora are essential for Machine Translation (MT) research and translation teaching. However, Arabic–English resources remain scarce, and existing datasets consist primarily of simple one-to-one mappings. In this paper, we present AlignAR, a generative sentence alignment method, and a new Arabic–English dataset comprising complex legal and literary texts. Our evaluation demonstrates that "Easy" datasets lack the discriminatory power to fully assess alignment methods. By reducing one-to-one mappings in our "Hard" subset, we exposed the limitations of traditional alignment methods. In contrast, LLM-based approaches demonstrated superior robustness, achieving an overall F1-score of 85.5%, a 9% improvement over previous methods. Our datasets and codes are open-sourced at https://github.com/XXX.
AlignAR: Generative Sentence Alignment for Arabic–English Parallel Corpora of Legal and Literary Texts
Baorong Huang1, Ali Asiri2thanks: Corresponding author: amaasiri@uqu.edu.sa, 1School of Foreign Languages, Huaihua University 2Al-lith University College, Umm al-Qura University
1 Introduction
Sentence alignment is a critical step in preparing parallel data for various Natural Language Processing (NLP) applications, particularly machine translation. It involves identifying corresponding sentences between two texts that are mutual translations. Given that manual alignment is resource-intensive and time-consuming, the field has witnessed significant development of automatic methods to address this fundamental challenge in computational linguistics.
Over the past three decades, sentence alignment research has progressed from early statistical and length-based models Brown et al. (1991); Gale and Church (1993) to heuristic, structural, and modern neural embedding-based approaches capable of handling diverse domains. Tools such as AIlign exemplify recent deep-learning innovations, combining multilingual sentence embeddings, anchor-point detection, and dynamic alignment algorithms. More recent work explores optimization-driven methods—such as Earth Mover’s Distance for biomedical alignment Luo et al. (2023) , dictionary-based extraction for Arabic–English comparable corpora Althobaiti (2022), and Particle Swarm Optimization for English corpora Shang and Li (2024). Despite these advances, research on sentence alignment using Large Language Models (LLMs) remains notably scarce, and current LLM-generated alignments often suffer from segmentation inconsistencies, many-to-many confusion, and context-loss errors.
Figure 1: Examples of 1-to-1, 1-to-many and many-to-many alignments between source and target sentences written in Arabic and English, respectively.
Our study aims to fill this gap by developing an Arabic–English alignment workflow that trains, evaluates, and systematically corrects the output of initial automated LLM alignment. Building on this, we introduce the Arabic-English Aligner, a bilingual annotation tool designed to create gold-standard corpora by enabling human operators to refine LLM-generated alignments through targeted operations such as merge, split, and exchange. This hybrid LLM-plus-human-validation model addresses deficiencies in current automated approaches and contributes a reliable framework for large-scale legal parallel corpus development.
Table 1: Statistics of the Arabic–English Parallel Corpus
Dataset #AR-SENT #AR-TKN #EN-SENT #EN-TKN SENT. (%) 1-1 (%)
Easy 1 153 4,405 206 7,783 74.27 78.15
Easy 2 202 6,437 257 11,438 78.60 77.78
Easy 3 161 4,726 198 8,471 81.31 81.76
Easy 4 202 4,893 233 8,766 87.73 85.20
Easy 5 174 5,770 199 10,650 87.44 87.21
Hard 1 93 3,880 202 6,318 46.03 28.77
Hard 2 100 3,998 194 6,823 51.55 40.26
Hard 3 93 3,111 175 5,849 53.14 45.78
Hard 4 92 3,068 203 5,800 45.32 40.70
Hard 5 101 2,759 223 6,906 45.29 31.18
[…]
Note: #AR-SENT and #AR-TKN indicate sentence count and token count in the Arabic text; #EN-SENT and #EN-TKN indicate sentence count and token count in the English text; 1-1 (%) means the frequency (percentage) of 1-to-1 alignments in each dataset. Both source and target texts have been tokenized with Stanza before counting tokens.
2 Existing methods
2.1 BleuAlign
Sennrich and Volk (2010) proposed a MT-based sentence alignment approach, commonly known as Bleualign, designed to address alignment challenges in noisy parallel data such as OCR-generated texts. Instead of relying solely on sentence length or lexical overlap, their method translates the source-language sentences using a machine translation system and then measures similarity between the translated sentences and target-language sentences using BLEU scores. High-scoring sentence pairs are treated as anchor points, while dynamic programming is employed to infer optimal alignments between anchors, allowing for one-to-one and one-to-many mappings. The approach also integrates traditional length-based heuristics as a fallback mechanism in difficult cases. Experimental results demonstrate that Bleualign outperforms classical length-based methods, particularly in noisy or structurally divergent corpora.
2.2 VecAlign
Thompson and Koehn (2019) introduced Vecalign, a sentence alignment algorithm that leverages multilingual sentence embeddings to efficiently align parallel documents. Vecalign represents sentences from both languages in a shared semantic vector space using pre-trained multilingual encoders and computes cosine similarity to identify semantically corresponding sentence pairs. To ensure scalability, the authors proposed a linear-time dynamic programming algorithm that aligns sentences monotonically while permitting sentence splits and merges. Unlike MT-based methods, Vecalign does not require translation at alignment time, significantly reducing computational cost. Evaluations across multiple language pairs show that Vecalign achieves higher recall and competitive precision compared to traditional aligners.
2.3 BertAlign
Liu and Zhu (2023) proposed Bertalign, a sentence alignment method that exploits contextualized representations generated by BERT-based multilingual transformers. In this approach, sentences are encoded using transformer embeddings that capture deep semantic and contextual information, enabling more precise similarity estimation than static word or sentence embeddings. Bertalign computes sentence-level similarity scores and applies dynamic programming to determine optimal alignments under monotonicity constraints, supporting one-to-one and one-to-many alignments. Experimental evaluations on parallel corpora show that Bertalign consistently outperforms earlier embedding-based and MT-based aligners, particularly in structurally divergent language pairs.
3 Datasets
The data used in this study are divided into two main categories based on the complexity of the sentence alignment task: hard-alignment data and easy-alignment data, as shown in Figure  1. the hard part consists of 5 documents, totaling 378 Arabic sentences and 774 English sentences, with the smallest source/target ratio around 0.45, which means that one Arabic sentence aligns to at least two English sentences on average. The easy part also consists of 5 documents manually aligned from Arabic laws and its English translations, totaling 892 source lines and 1093 target lines, with the smallest source/target ratio around 0.74, much higher than the hard part.
Figure 2: Our alignment procedure consists of three steps: (1) sentence indexing, i.e., explicitly indexing the lines with sequence number; (2) LLM-based translation mapping, i.e., requesting the selected LLM to select the translations from the target lines; (3) ladder building, extracting the index number from the formatted JSON output and obtaining ladder files for evaluation.
The data of hard-alignment subset are drawn from classical and literary Arabic texts, which present significant challenges due to archaic language, stylistic density, and non-literal translation strategies. In this category, the study utilizes sentence-aligned parallel data from classical Arabic books and their English translations made available through the open-source platform Rasaif. In particular, the philosophical narrative Ḥayy ibn Yaqẓān was extracted and manually verified at the sentence level, providing a challenging testbed for alignment due to its complex syntax, metaphorical language, and interpretive translations. Additional literary data were obtained from short stories included in Hassan Ghazala’s A Textbook of Literary Translation (2013), which feature rich stylistic variation and frequent sentence restructuring across languages.
In contrast, the easy-alignment subset comprises modern legal texts characterized by high structural parallelism and terminological consistency. These data were collected from the Saudi National Center for Archives and Records (NCAR), which provides official Saudi laws alongside their English translations. Several statutes—including the Companies Law, Anti-Bribery Law, and Anti-Commercial Concealment Law—were manually collected, cleaned, and aligned at the sentence level in both languages. Due to their formulaic drafting style, controlled vocabulary, and close translation correspondence, these legal texts represent low-complexity alignment data. The inclusion of both challenging literary material and structurally aligned legal texts enables the model to learn alignment patterns across a wide spectrum of linguistic and translational complexity, strengthening its generalizability and robustness for English–Arabic sentence alignment tasks.
Figure 3: LLMAligner
The alignment are manually done with our open-source Arabic-English Aligner named LLMAligner, as shown in Figure 3. The tool supports the manual adjustment of the mappings, including Merge, Split, and Exchange, and exporting the results. The source code of the tool is available at GitHub Repository.
4 Method
In this study, we frame document alignment as a zero-shot inference task, prompting the models to identify correspondences based on semantic equivalence rather than lexical overlap. We propose a novel sentence alignment that considered alignment as a translation mapping task, and request the LLM to identify the translations of the source text, rather than directly performing the alignment.
4.1 Alignment Procedure
We propose a three-step alignment procedure. In the first step, we intuitively add a sentence index (sequence number) as $0\ s_{0},\ \dots,\ n\ s_{m}$] and $0\ t_{0},\ \dots,\ n\ t_{n}$] to each source line and each target line to reduce the hallucinations produced by an LLM. To further reduce the inadvertent addition or deletion of the words by an LLM, we change the parallel aligning task into a translation mapping tasks, requesting the LLM to pick out the corresponding translations of the source lines from the target lines and output the entire mappings in a JSON format. In the final step, we extract the sentence index to obtain a ladder file that only contains line number mappings for evaluation, as shown in Figure 2.
4.2 Evaluation Metrics
We used the strict precision, strict recall and strict F1 as our evaluation metrics. Strict true positive, as defined by Sennrich and Volk (2010), means the alignment hypothesis should be identical to the reference alignment. We adopt the strict metrics because the alignment scores produced by VecAlign and BleuAlign had already exceeded 90% and the actual usage of the alignments in translation studies or MT training requires the precise and exact mappings between the source lines and target lines. In this case, the overlapping alignment as lax measures allowed will be considered incorrect. We takes the Micro-Averaged Metrics for computing the overall scores for the easy and hard parts.
5 Experiments & Results
5.1 Experimental settings
We evaluate the alignment accuracy on the Arabic-English datasets using our proposed LLM-based alignment methods, in comparison with the previous methods, including BertAlign, BleuAlign and VecAlign. Our experiments are conducted on a computer with Intel Ultra 7 265KF, 64GB RAM and NVIDIA RTX 5070 Ti. Two LLMs are used in the experiments, including GPT-5.1-mini and Gemini-2.5-flash. GPT is accessed via official API while Gemini is accessed via OpenRouter.
5.2 Experiments on legal text (easy part)
An analysis of the experimental results in Table 2 suggests that while traditional Transformer-based models like BERT maintain high performance, the "Easy" subsets of the dataset may lack the discriminatory power required to fully evaluate the capabilities of aligning methods. As shown by the "Overall" metrics, Gemini achieves a near-perfect $F_{1}$ score of 0.993; however, the consistently high scores across all methodologies, all exceeding the 0.90 threshold, indicate that these specific benchmarks have reached saturation. Consequently, more challenging data are required to reveal the latent performance gaps.
Table 2: Comparison across different methods on easy document alignments
Dataset Metric Baseline approaches Our approach
BERT BLEU VEC GPT Gemini
Easy 1 P 0.936 0.966 0.961 0.987 0.987
R 0.961 0.947 0.961 0.993 0.993
F1 0.948 0.957 0.961 0.990 0.990
Easy 2 P 0.985 0.919 0.985 0.990 0.980
R 0.980 0.914 0.985 0.995 0.990
F1 0.982 0.916 0.985 0.993 0.985
Easy 3 P 0.981 0.925 0.975 1.000 1.000
R 0.981 0.925 0.956 1.000 1.000
F1 0.981 0.925 0.965 1.000 1.000
Easy 4 P 1.000 0.944 0.970 0.980 0.990
R 1.000 0.954 0.965 0.990 0.995
F1 1.000 0.949 0.967 0.985 0.992
Easy 5 P 1.000 0.965 0.994 0.989 1.000
R 1.000 0.965 0.988 0.994 1.000
F1 1.000 0.965 0.991 0.991 1.000
Overall P 0.982 0.943 0.977 0.989 0.991
R 0.985 0.943 0.972 0.994 0.995
F1 0.984 0.942 0.974 0.992 0.993
BERT means BertAlign, BLEU means BlueAlign and VEC means VecAlign. Bold indicates the best-performing method.
5.3 Experiments on literary text (hard part)
While the preliminary evaluations on simplified datasets suggested a performance parity among various models, the results on the challenging document alignment task (Table 3) provide a much clearer differentiation of model capabilities. As the task complexity increases, the "ceiling effect" observed in the easy datasets disappears, revealing significant discrepancies in model robustness. Gemini emerges as the superior approach, achieving an overall $F_{1}$ score of 0.855, which represents a substantial margin over the BERT baseline (0.726) and the GPT-based approach (0.767). The sharp decline in performance across all baseline methods when moving from "Easy" to "Hard" datasets underscores the limitations of traditional semantic matching. For instance, BLEU’s overall $F_{1}$ plummeted from 0.942 to 0.477, and BERT’s overall $F_{1}$ dropped from 0.984 to 0.726. More importantly, the divergence between GPT and Gemini on the hard datasets—most notably in Hard 5, where Gemini outperforms GPT by over 12 percentage points in $F_{1}$—indicates that the selection of LLMs also exterts considerable influence on the alignment quality.
Table 3: Comparison across different methods on hard document alignments
Dataset Metric Baseline approaches Our approach
BERT BLEU VEC GPT Gemini
Hard 1 P 0.667 0.507 0.670 0.684 0.811
R 0.757 0.479 0.735 0.783 0.880
F1 0.726 0.493 0.701 0.730 0.844
Hard 2 P 0.724 0.563 0.791 0.718 0.871
R 0.807 0.519 0.818 0.841 0.920
F1 0.763 0.540 0.804 0.775 0.895
Hard 3 P 0.606 0.410 0.645 0.612 0.688
R 0.648 0.410 0.682 0.682 0.750
F1 0.626 0.410 0.663 0.645 0.717
Hard 4 P 0.778 0.618 0.826 0.884 0.903
R 0.865 0.547 0.854 0.944 0.944
F1 0.819 0.580 0.840 0.913 0.923
Hard 5 P 0.699 0.340 0.780 0.748 0.880
R 0.742 0.310 0.804 0.794 0.907
F1 0.720 0.324 0.792 0.770 0.893
Overall P 0.696 0.494 0.743 0.729 0.831
R 0.757 0.462 0.780 0.809 0.831
F1 0.726 0.477 0.761 0.767 0.855
6 Conclusions
In this paper, we introduced a generative sentence alignment method, and a new Arabic–English parallel dataset focused on legal and literary texts. Our findings reveal that standard "Easy" datasets lack the discriminatory power to fully evaluate model capabilities. By reducing one-to-one mappings in our "hard" subset, we exposed the limitations of previous method. In contrast, the LLM-based methods demonstrated superior robustness, maintaining high precision and recall where other methods degraded.
Limitations
The proposed LLM-based document aligning methods can effectively and accurately align the low-resource Arabic text to English text. However, the required time for the alignment depends on the processing time of the LLMs. In addition, the format of the response is also essential for the successful parsing and construction of the alignment mappings. Furthermore, the length of the document is also constrained by the context window of the large language models. In our experiments, one simple prompt is used, and the effects of advanced prompt methods, such as few-shots, chain-of-reason (COT), self-reflections or agentic design, are not explored.
Acknowledgments
The authors extend their appreciation to Umm Al-Qura University, Saudi Arabia for funding this research work through grant number: 26UQU4350258GSSR01.
Funding statement
This research was funded by Umm Al-Qura University, Saudi Arabia under grant number 26UQU4350258GSSR01.